# Cluster Catch Digraphs with the Nearest Neighbor Distance

### v1

@@ 6 Real and Complex Data sets
Consider the D31 data set, KS-CCDs and K-means++ detect all 31 clusters, deliver the highest ARIs of over 0.9, outperforming other methods. RK-CCDs and UN-CCDs also perform well, with ARI scores of 0.814 and 0.799, respectively. They detect 30 and 31 clusters, respectively, outperforming all other methods except KS-CCDs and K-means++.
@@
In summary, KS-CCDs demonstrate the best overall performance, obtaining the highest Sil and competitive ARI across almost all data sets. However, identifying the best density parameter $\delta$ for the KS-based statistic is computationally expensive. Even when the number of clusters is provided as an input parameter for K-means++ and spectral clustering, RK-CCDs and UN-CCDs maintain comparable performance, making them competitive clustering methods. DBSCAN and Louvain clustering method yield slightly worse results than the five methods mentioned above. MST consistently produces negative ARI or Silhouette scores, indicating the worst performance among the evaluated methods.
@@ 7 Summary and Conclusion
UN-CCDs address the limitation of RK-CCDs in high-dimensional clustering, attributes to several enhancements discussed in Section 4.2. Through extensive Monte Carlo simulation in Section 5, we demonstrate that UN-CCDs are, in general, comparable to RK-CCDs in low-dimensional space, and deliver superior performance when the dimensionality $d\geq 10$, This is evidenced by the substantially higher Adjust Rand Index and the success rate in identifying the correct number of clusters. Moreover, as shown by the real and complex data analysis in Section 6, UN-CCDs maintain comparable performance to RK-CCDs and state-of-the-art methods such as spectral clustering and K-means++ (which require the number of clusters as an input parameter). These findings establish UN-CCDs as a robust and versatile clustering method, suitable for a wide range of data dimensionality and complexity.

### v2

@@ 5 Real Data Examples
Consider the “D31” dataset, KS-CCDs and $K$-means++ detect all 31 clusters, deliver the highest ARI values of over 0.9, outperforming other methods. RK-CCDs and UN-CCDs also perform well, with ARI values of 0.814 and 0.799, respectively. They detect 30 and 31 clusters, respectively, outperforming all other methods except KS-CCDs and $K$-means++.
Summary of the Experiment:
@@
Overall, KS-CCDs demonstrate the best performance, obtaining the highest Sil and competitive ARI values across almost all datasets. However, identifying the best density parameter $\delta$ for the KS-type statistic is computationally expensive. Even when the number of clusters is provided as an input parameter for $K$-means++ and spectral clustering, RK-CCDs and UN-CCDs maintain comparable performance, with UN-CCDs slightly better making them competitive clustering methods. DBSCAN and Louvain clustering methods yield slightly worse results than RK-CCDs. MST consistently produces negative ARI or Silhouette values, indicating the worst performance among the evaluated methods, since it is not robust to noise/outliers and the varying intensities of clusters.
@@ 6 Summary and Conclusion
UN-CCDs address the limitation of RK-CCDs in high-dimensional clustering, attributes to several enhancements discussed in Section 3.2. Through extensive Monte Carlo simulation in Section 4, we demonstrate that UN-CCDs are, in general, comparable to RK-CCDs in low-dimensional space, and deliver superior performance when the dimensionality $d\geq 10$, This is evidenced by the substantially higher Adjust Rand Index and the success rate in identifying the correct number of clusters. Moreover, as shown by the real and complex data analysis in Section 5, UN-CCDs maintain comparable performance to RK-CCDs and state-of-the-art methods such as spectral clustering and K-means++ (which require the number of clusters as an input parameter). These findings establish UN-CCDs as a robust and versatile clustering method, suitable for a wide range of data dimensionality and complexity. However, the UN-CCD method comes with some limitations: (1) It has high time complexity of $O(n^{3})$. (2) It suffers reduced effectiveness in curtain in specific scenarios, e.g., when the dataset is consisted of non-uniform clusters with low cluster intensities and close proximity. (3) The method currently requires manual tuning of the significance level ($\alpha$) for its spatial randomness test.
Future Work:
Lastly, we enumerate some interesting future work directions for UN-CCDs, including addressing the limitations of UN-CCDs and applying them to develop other clustering methods for different applications.
  • (1)
    Extending UN-CCDs for Overlapping Clusters, developing methods similar to Fuzzy clustering.
  • (2)
    Applying CCD-Based Clustering in Semi-Supervised Learning for partially labeled data.
  • (3)
    Adaptive neighborhood or radius selection based on local density.
  • (4)
    Developing automated process for tuning the significance levels ($\alpha$) for the MC-SRT with NND, satisfying the need of different datasets.
  • (5)
    The time complexity of the UN-CCD method is $O(n^{3})$. Although it is faster than RK-CCDs, there is a substantial room for improvement.